OpenStudy (anonymous):

if angle M and angle N are supplementary angles, and angle M and angle P are complementary angles, what is the relationship between angle N and angle P.

OpenStudy (anonymous):

M + N = 180 M + P = 90 M + N = 180 M = 180 - N 180 - N + P = 90 N = 90 + P

OpenStudy (2bornot2b):

N-P=90

OpenStudy (mertsj):

n-p=m

OpenStudy (anonymous):

 n and p must also be complementary